# Optimal Average Disk-Inspection via Fermat’s Principle

### v2

@@ 3.2 Disk-Inspection via Discretization (and Nonlinear Programming)
As $k\to\infty$, $(\theta,k)$-inspective curves approximate $\theta$-inspective curves. In particular, scaling a $(\theta,k)$-inspective curve by a factor $1+O(1/k)$ produces a $\theta$-inspective curve, which is one interpretation of the upper bounds derived in [16].
@@
It was shown in [16] that if $t_{i}=\Omega(1/k)$ for all $i$, then the chain is also a $\theta$-inspective curve. As $k\to\infty$, $C_{\theta,k}(t)$ converges to the cost $s(\theta)$ of the $\theta$-ADI defined in Theorem 3.3. The expression $C_{\theta,k}(t)$ is non-convex in the variables $t$ and $\theta$. Under the feasibility constraints $t_{i}=\Omega(1/k)$, it serves as the objective of a NonLinear Program characterizing the cost of any solution to ADI. This is precisely the NLP formulation used in [16] to derive upper bounds. In what follows, we analyze how to reduce the number of variables to that NLP down to one variable and examine the limiting behavior as $k\to\infty$, resulting in the promised SPOCP.
@@ 4 Fermat’s Principle Solves $(\theta,k)$-ADI
In this section we characterize the optimal trajectory to $(\theta,k)$-ADI and its cost via a recursion based on Fermat’s Principle. This marks the beginning of our new contributions.
@@ 4.2 Optimal Solution to $(\theta,k)$-ADI via Recursion
Given $\theta\in[0,\pi/2)$ and $k\geq 5$, suppose the optimal trajectory to $(\theta,k)$-ADI is identified by $A_{i}=L_{i}(t_{i})$ with $t_{i}\geq\tan\mkern-2.0mu\left(\tfrac{\alpha}{2}\right)$ (so it does not intersect the unit disk). [^2] Note that $\alpha\rightarrow 0$, as $k\rightarrow\infty$, so if $t_{i}\geq\epsilon$ for some constant $\epsilon>0$, the induced trajectory does not intersect the disk, for large values of $k$. Then the trajectory and its cost are characterized by the recursions
@@
Consider triangle $A_{i}A_{i-1}R_{i}$, where $R_{i}$ is the intersection of $L_{i}$ and $L_{i-1}$. Since the lemma assumes $t_{j}\geq\tan\mkern-2.0mu\left(\tfrac{\alpha}{2}\right)$ for all $j$, in particular $t_{i-1}\geq\tan\mkern-2.0mu\left(\tfrac{\alpha}{2}\right)$, hence $\|R_{i}A_{i-1}\|=t_{i-1}-\tan\mkern-2.0mu\left(\tfrac{\alpha}{2}\right)\geq 0$, so triangle $A_{i}A_{i-1}R_{i}$ is well defined. The angle at $A_{i}$ is $x_{i}$, the angle at $A_{i-1}$ is $\pi-y_{i-1}$, and the angle at $R_{i}$ equals $\alpha$ (from quadrilateral $OP_{i}R_{i}P_{i-1}$). Thus $x_{i}+(\pi-y_{i-1})+\alpha=\pi$, which gives (7).
@@
which, by Lemma 3.6, is proportional to $C_{\theta,k}(t)$ by a factor independent of $t$. Hence, by Lemma 4.2, refraction holds at each interface. At the interface between $M_{i}$ ($s_{i}=1/i$) and $M_{i+1}$ ($s_{i+1}=1/(i+1)$), the incidence and refraction angles with respect to the normal are $\alpha_{1}=\tfrac{\pi}{2}-x_{i}$ and $\alpha_{2}=\tfrac{\pi}{2}-y_{i}$. Therefore, by Snell’s Law (Theorem 4.1), we have $\tfrac{\sin\mkern-2.0mu\left(\alpha_{1}\right)}{\sin\mkern-2.0mu\left(\alpha_{2}\right)}=\tfrac{s_{i}}{s_{i+1}}=\tfrac{i+1}{i}.$ Since $\sin\mkern-2.0mu\left(\tfrac{\pi}{2}-z\right)=\cos\mkern-2.0mu\left(z\right)$, this becomes $\tfrac{\cos\mkern-2.0mu\left(x_{i}\right)}{\cos\mkern-2.0mu\left(y_{i}\right)}=\tfrac{i+1}{i}$, which together with (7) yields (8) .
@@ 5 Single-Parameter Reformulation and Continuum Limit
From this setup we obtain the continuum limit as $n\to\infty$, yielding the ODE system $\textsc{Sys}(\tau_{0})$ of Definition 2.5. The corresponding trajectory $\mathcal{T}$ begins at $(1,-\tau_{0})$, remains outside the unit disk, and intersects the line $x=1$ in the first quadrant at some point $\mathcal{T}(\xi)$, where $\xi$ is the deployment parameter. We now state a sequence of technical results, while we defer their proofs to Appendix C.
@@
Let $\phi_{i}$ and $L_{i}(t)$ be defined as in (4) and (5), and let $A_{i}=L_{i}(t_{i})$, where $(t_{i})$ is defined by (9) with $t_{0}=\tau_{0}$. Then the polygonal trajectory through $(A_{i})_{i=0}^{n}$ converges, as $n\to\infty$, to the curve $\mathcal{T}:[0,1]\to\mathbb{R}^{2}$ of Definition 2.5, with $\mathcal{T}(0)=A_{0}$.
@@ 5.2 Bounds on the Optimal Deployment Angle
For this reason, we must restrict the range $0,\pi/2)$ of deployment angles so as, first, to exclude angles for which the optimal trajectories touch the disk (small values of $\theta$), and second, to exclude angles for which the associated SPOCP becomes numerically unstable (large values of $\theta$). These restrictions are established through a combination of analytic arguments and numerical evaluations, the latter of which are justified in Section 5.4. As we show in the next section, the remaining range of deployment angles contains the minimizer, which we then compute.
To resume, in this section we show a refinement of (3), as follows.
@@
Set $\mathcal{O}=\inf_{\theta\in[0,\pi/2)}B_{\theta}(s)$. The main contribution to [16] is an upper bound of $3.5509015$ to the optimal cost to ADI, that is $\mathcal{O}\leq 3.5509015$. Next we show that $B_{\theta}(s)>\mathcal{O}$ when $\theta>1.148$.
@@
The inspection cost for points in $I_{1}$ is 0. For points in $I_{2}$, the cost is at least the time required to inspect $P_{\pi}$, namely $2$. For points in $I_{3}$, the cost is at least the time required to inspect $P_{3\pi/2}$. In this case we employ the provably optimal trajectory from [16], established for the worst-case problem. Since all preceding points must already have been inspected by moving counterclockwise around the disk, $P_{3\pi/2}$ cannot be inspected earlier than $\tan\mkern-2.0mu\left(\theta\right)+(\pi-2\theta)+1,$ where $\tan\mkern-2.0mu\left(\theta\right)$ is the tangent length to the disk, $\pi-2\theta$ is the circular arc length, and the additional $1$ is the final straight segment needed to complete visibility without intersecting the disk. Overall, this shows that
@@
Consider the optimal trajectory for the partial problem $\theta$-ADI which inspects all points $\{P_{\phi}\}_{\phi\in[2\theta,2\pi]}$, each with inspection time $I(P_{\phi})$. It follows that such a trajectory is also feasible to $(\theta,k)$-ADI problem that inspects $k$ equidistant points in the same arc. Note that any point $P_{\phi}$ in interval $\phi\in[\phi_{i},\phi_{i-1}]$ is inspected, by the triangle inequality, in time at least $\sum_{j=i+1}^{k}\|A_{j}A_{j-1}\|.$ It follows that for the cost to the continuous problem we have
@@
In this section, we show that the premise of Theorem 2.8 is satisfied, namely that the optimal inspective curve does not touch the unit disk. Consequently, this allows us to optimize expression (1) and thus conclude with the proof of Theorem 2.4. We rely on the theoretical foundations established previously, and, as is necessary, we make use of further numerical calculations implemented in the Julia programming language [11]. Many of the arguments below are based on numerical comparisons. In Section 5.4 we provide the justifications that these computations are robust and consistent with the accuracy promised in the main theorem.
@@
Therefore, in order to determine the optimal solution it remains to minimize the cost expression of Theorem 2.8 over the admissible trajectories. The ODE formulation shows that each trajectory is uniquely determined by the initial condition $\tau(0)=\tau_{0}$, so that the cost becomes a function of this single parameter. Hence the problem reduces to $\textsc{SPOCP}(\tau_{0})$, namely the problem of minimizing (1) over $\tau_{0}\in[1.64697,1.6525]$. Figure 6 summarizes the numerical evaluation over increasingly refined intervals of 2000 grid points each. The minimum is sandwiched between $3.5492598$ and $3.54925986$. For $\tau_{0}=1.6469768608776936$ (exact value) we obtain $\xi=0.8119098734258519\ldots$, $\theta=0.5909025598581181\ldots$, and a corresponding inspective curve with minimum distance at least $0.0302318\ldots$ to the disk boundary (corresponding to $\tau_{(}\chi)\approx 0.24774522\ldots$ for some $\chi$), and reported cost approximately $3.5492595860809693\ldots$. Numerical accuracy of this computation is discussed in Section 5.4. ∎
@@ 5.4 Numerical Methods and Accuracy Guarantees
We used the Julia programming language [11] for all computations.
For the convex nonlinear program that lower bounds $s(\theta)$ in Lemma 5.9, we model it with JuMP [19] and solve it using the interior point method Ipopt [14, 41]. The program is convex since the feasible set $\{t_{i}\geq 0\}$ is convex and the objective is a nonnegative conical combination of norms of affine maps. Hence any KKT point is globally optimal. In practice, the solver returns primal and dual feasible solutions with residuals below $10^{-14}$ and objective values stable to at least $10^{-9}$. Because convexity guarantees global optimality, these certificates validate the solutions and support the digits reported in our lower bounds.
For the ODE system of Definition 2.5, used in Section 5.3 to establish feasibility of trajectories and to evaluate the cost functional, we proceed as follows. The equation for $\psi$ has a singularity at $x=0$, so we begin the integration at $x_{0}=10^{-6}$ using the asymptotic expansion $\psi(x_{0})=\pi/2-\pi x_{0}+(\pi^{2}/2)x_{0}^{2}$, whose truncation error is $O(x_{0}^{3})\approx 10^{-18}$ and therefore negligible compared with the solver tolerances. We integrate $\psi$ on $x_{0},1]$ with the adaptive stiff solver Rodas5() from the DifferentialEquations.jl library [40], setting absolute and relative tolerances to $10^{-12}$, and then integrate $\tau$ on the same interval using the dense output of $\psi$ in the right-hand side.
The deployment parameter $\xi$ is obtained in two stages. A uniform grid of 10,00010{,}000 points provides a coarse $\xi_{\text{approx}}$ as the last feasible point according to a geometric predicate that tolerates floating-point noise at the level $10^{-12}$. This value is refined by bisection to absolute tolerance $10^{-8}$. To further confirm stability, we repeat the bisection at tolerance $5\cdot 10^{-10}$ and recompute the objective, reporting the discrepancy between the two runs, which is consistently negligible. In addition, the minimum of $\tau(\cdot)$ on $x_{0},\xi_{\text{approx}}]$ is computed by Brent’s method [13] with tolerance $10^{-10}$ to certify that all trajectories remain uniformly outside the disk.
The integral of Lemma 5.4 is evaluated using the adaptive Gauss-Kronrod quadrature routine quadgk from QuadGK.jl [32], with relative tolerance $10^{-12}$ and absolute tolerance $10^{-14}$. Substituting $(\xi,\theta)$ into the expression of Theorem 2.8 then yields the values reported in Section 5.3.

### v3

@@ 3.2 Disk-Inspection via Discretization (and Nonlinear Programming)
We assume $k\geq 5$ so that the angular spacing between consecutive discretization angles is strictly less than $\pi/2$; this condition is required to ensure that feasibility of the discretized formulation implies feasibility of the corresponding continuous inspection trajectory, as already used in [conley2025multi].
@@
As $k\to\infty$, $(\theta,k)$-inspective curves approximate $\theta$-inspective curves. In particular, scaling a $(\theta,k)$-inspective curve by a factor $1+O(1/k)$ produces a $\theta$-inspective curve, which is one interpretation of the upper bounds derived in [conley2025multi].
@@
It was shown in [conley2025multi] that if $t_{i}=\Omega(1/k)$ for all $i$, then the polygonal chain induces a $\theta$-inspective curve for the corresponding continuous instance. Under this feasibility requirement, the discrete average $C_{\theta,k}(t)$ is a Riemann-sum approximation of the continuous average inspection cost associated with the induced $\theta$-inspective curve, since inspection times vary continuously with the endpoint angle at scale $O(1/k)$. In particular, optimizing $C_{\theta,k}(t)$ over feasible discretizations yields asymptotically tight upper bounds on the continuous $\theta$-ADI cost, as used in [conley2025multi].
The expression $C_{\theta,k}(t)$ is non-convex in the variables $t$ and $\theta$. Under the feasibility constraints $t_{i}=\Omega(1/k)$, it serves as the objective of a nonlinear program characterizing the cost of discretized solutions to ADI. In what follows, we show how this program can be reduced to a single-parameter optimization problem and analyze its limiting behavior as $k\to\infty$, resulting in the promised SPOCP.
@@ 4 Fermat’s Principle Solves $(\theta,k)$-ADI
In this section we characterize optimal trajectories to $(\theta,k)$-ADI and their cost via a recursion based on Fermat’s Principle. This marks the beginning of our new contributions.
@@ 4.2 Optimal Solution to $(\theta,k)$-ADI via Recursion
Given $\theta\in[0,\pi/2)$ and $k\geq 5$, suppose an optimal trajectory to $(\theta,k)$-ADI is identified by $A_{i}=L_{i}(t_{i})$ with $t_{i}\geq\tan\mkern-2.0mu\left(\tfrac{\alpha}{2}\right)$ (so it does not intersect the unit disk). [^2] The condition $t_{i}\geq\tan\mkern-2.0mu\left(\tfrac{\alpha}{2}\right)$ is the standard geometric feasibility requirement in the $(\theta,k)$-inspection framework of [conley2025multi], ensuring that the discrete trajectory does not intersect the unit disk. We note that in the parameter ranges relevant to our later analysis and numerical evaluations (with $k$ large), the optimized values of $t_{i}$ are observed to be bounded well away from 0, so this condition is comfortably satisfied in practice. Then the trajectory and its cost are characterized by the recursions
@@
Figure 4: Close-up detail of Figure 3 showing point $R_{i}$ obtained by the intersection of the lines tangent on the disc at points $P_{i},P_{i-1}$.
Consider triangle $A_{i}A_{i-1}R_{i}$, where $R_{i}$ is the intersection of the extensions of halflines $L_{i}$ and $L_{i-1}$ to full lines, see Figure 4. By the definition of $t_{i}$, we have that $\|P_{i-1}A_{i-1}\|=t_{i-1}$.
Moreover, $R_{i}$ is the intersection of two tangents at points $P_{i-1},P_{i}$ whose central angle is $\alpha$. Since $\ell_{i-1}\perp OP_{i-1}$ and $\ell_{i}\perp OP_{i}$, the line $OR_{i}$ bisects $\angle P_{i-1}OP_{i}$, hence $\angle P_{i-1}OR_{i}=\alpha/2$. In the right triangle $OP_{i-1}R_{i}$ we obtain $\|P_{i-1}R_{i}\|=\tan\mkern-2.0mu\left(\tfrac{\alpha}{2}\right)$. Since the lemma assumes $t_{j}\geq\tan\mkern-2.0mu\left(\tfrac{\alpha}{2}\right)$ for all $j$, in particular $t_{i-1}\geq\tan\mkern-2.0mu\left(\tfrac{\alpha}{2}\right)$, we have that $\|R_{i}A_{i-1}\|=t_{i-1}-\tan\mkern-2.0mu\left(\tfrac{\alpha}{2}\right)\geq 0$. The angle at $A_{i}$ is $x_{i}$, the angle at $A_{i-1}$ is $\pi-y_{i-1}$, and the angle at $R_{i}$ equals $\alpha$ (from quadrilateral $OP_{i}R_{i}P_{i-1}$). Thus $x_{i}+(\pi-y_{i-1})+\alpha=\pi$, which gives (7).
@@
which, by Lemma 3.6, is proportional to $C_{\theta,k}(t)$ by a factor independent of $t$. Hence, by Lemma 4.2, refraction holds at each interface. Indeed, fix $i\in\{1,\ldots,k-1\}$ and consider the interface line associated with $L_{i}$ separating the two media $M_{i}$ and $M_{i+1}$. If we keep the points $A_{i-1}\in L_{i-1}$ and $A_{i+1}\in L_{i+1}$ fixed, then any feasible trajectory induces a unique crossing point $A_{i}\in L_{i}$, and the contribution of the two adjacent segments to the travel time equals
$i\|A_{i}-A_{i-1}\|+(i+1)\|A_{i+1}-A_{i}\|.$
All other terms in the total travel time are independent of the choice of $A_{i}$. Since the trajectory is optimal, the chosen point $A_{i}$ must minimize the above expression over $A_{i}\in L_{i}$. Therefore Lemma 4.2 applies at $L_{i}$, and the segment directions satisfy Snell’s law between $M_{i}$ and $M_{i+1}$. If refraction failed at some interface, replacing $A_{i}$ by the minimizer given by Lemma 4.2 would strictly decrease the total travel time, contradicting optimality.
At the interface between $M_{i}$ ($s_{i}=1/i$) and $M_{i+1}$ ($s_{i+1}=1/(i+1)$), the incidence and refraction angles with respect to the normal are $\alpha_{1}=\tfrac{\pi}{2}-x_{i}$ and $\alpha_{2}=\tfrac{\pi}{2}-y_{i}$. Therefore, by Snell’s Law (Theorem 4.1), we have $\tfrac{\sin\mkern-2.0mu\left(\alpha_{1}\right)}{\sin\mkern-2.0mu\left(\alpha_{2}\right)}=\tfrac{s_{i}}{s_{i+1}}=\tfrac{i+1}{i}.$ Since $\sin\mkern-2.0mu\left(\tfrac{\pi}{2}-z\right)=\cos\mkern-2.0mu\left(z\right)$, this becomes $\tfrac{\cos\mkern-2.0mu\left(x_{i}\right)}{\cos\mkern-2.0mu\left(y_{i}\right)}=\tfrac{i+1}{i}$, which together with (7) yields (8) .
@@ 5 Single-Parameter Reformulation and Continuum Limit
This setup leads to the following continuum limit as $n\to\infty$, namely the ODE system $\textsc{Sys}(\tau_{0})$ of Definition 2.5. The corresponding trajectory $\mathcal{T}$ begins at $(1,-\tau_{0})$, remains outside the unit disk, and intersects the line $x=1$ in the first quadrant at some point $\mathcal{T}(\xi)$, where $\xi$ is the deployment parameter. We now state a sequence of technical results establishing this limit. We defer their proofs to Appendix C.
@@
Let $\phi_{i}$ and $L_{i}(t)$ be defined as in (4) and (5), and let $A_{i}=L_{i}(t_{i})$, where $(t_{i})$ is defined by (9) with $t_{0}=\tau_{0}$. Assume that $\tau_{0}$ is inspection-feasible (Definition 2.7), and let $\xi=\xi(\tau_{0})$ be its deployment parameter, that is, the smallest $\xi\in(0,1]$ such that $\mathcal{T}_{1}(\xi)=1$ for $\mathcal{T}=\mathcal{T}^{\tau_{0}}$ of Definition 2.6. Let $\widetilde{\mathcal{T}}_{n}$ be the polygonal path obtained by linear interpolation through the points $(A_{i})_{i=0}^{n}$ with parameter values $h_{i}=i/n$.
Then $\widetilde{\mathcal{T}}_{n}$ converges to $\mathcal{T}$ on $0,\xi)$, in the sense that for every $0<\delta\leq\xi^{\prime}<\xi$,
$\sup_{x\in[\delta,\xi^{\prime}]}\lVert\widetilde{\mathcal{T}}_{n}(x)-\mathcal{T}(x)\rVert\to 0\qquad\text{as }n\to\infty.$
Moreover, $\mathcal{T}(0)=A_{0}$.
@@ 5.2 Bounds on the Optimal Deployment Angle
For this reason, we restrict the range $0,\pi/2)$ of deployment angles. Small values of $\theta$ are excluded because the corresponding optimal trajectories touch the disk. Large values of $\theta$ are excluded for a different reason: as $\theta\to\pi/2$, the deployment point moves arbitrarily far from the origin, and we show in the next section that such angles cannot be optimal. Consequently, these angles need not be considered in the optimization. We then search numerically for the minimizing deployment angle over the remaining range. Restricting to this range also avoids numerical instability in the associated SPOCP, which would otherwise arise from the large scales involved. The numerical procedures used in this restricted regime are justified in Section 5.4.
To summarize, in this section we show a refinement of (3), as follows.
@@
Set $\mathcal{O}=\inf_{\theta\in[0,\pi/2)}B_{\theta}(s)$. The main contribution to [conley2025multi] is an upper bound of $3.5509015$ to the optimal cost to ADI, that is $\mathcal{O}\leq 3.5509015$. Next we show that $B_{\theta}(s)>\mathcal{O}$ when $\theta>1.148$.
@@
The inspection cost for points in $I_{1}$ is 0. For points in $I_{2}$, the cost is at least the time required to inspect $P_{\pi}$, namely $2$. For points in $I_{3}$, the cost is at least the time required to inspect $P_{3\pi/2}$. In this case we employ the provably optimal trajectory from [conley2025multi], established for the worst-case problem. Since all preceding points must already have been inspected by moving counterclockwise around the disk, $P_{3\pi/2}$ cannot be inspected earlier than $\tan\mkern-2.0mu\left(\theta\right)+(\pi-2\theta)+1,$ where $\tan\mkern-2.0mu\left(\theta\right)$ is the tangent length to the disk, $\pi-2\theta$ is the circular arc length, and the additional $1$ is the final straight segment needed to complete visibility without intersecting the disk. Overall, this shows that
@@
Consider an optimal trajectory for the partial problem $\theta$-ADI which inspects all points $\{P_{\phi}\}_{\phi\in[2\theta,2\pi]}$, each with inspection time $I(P_{\phi})$. It follows that such a trajectory is also feasible to $(\theta,k)$-ADI problem that inspects $k$ equidistant points in the same arc. Note that any point $P_{\phi}$ in interval $\phi\in[\phi_{i},\phi_{i-1}]$ is inspected, by the triangle inequality, in time at least $\sum_{j=i+1}^{k}\|A_{j}A_{j-1}\|.$ It follows that for the cost to the continuous problem we have
@@
In this section, we show that the premise of Theorem 2.8 is satisfied, namely that the optimal inspective curve does not touch the unit disk. Consequently, this allows us to optimize expression (1) and thus conclude with the proof of Theorem 2.4. We rely on the theoretical foundations established previously, and, as is necessary, we make use of further numerical calculations implemented in the Julia programming language [bezanson2017julia]. Many of the arguments below are based on numerical comparisons. In Section 5.4 we provide the justifications that these computations are robust and consistent with the accuracy promised in the main theorem.
@@
Therefore, in order to determine the optimal solution it remains to minimize the cost expression of Theorem 2.8 over the admissible trajectories. The ODE formulation shows that each trajectory is uniquely determined by the initial condition $\tau(0)=\tau_{0}$, so that the cost becomes a function of this single parameter. Hence the problem reduces to $\textsc{SPOCP}(\tau_{0})$, namely the problem of minimizing (1) over $\tau_{0}\in[1.64697,1.6525]$.
Since the admissible family is one-dimensional, we evaluate this function numerically and refine the search interval around the minimizer. Figure 7 summarizes this coarse-to-fine evaluation over increasingly refined intervals of 2000 grid points each. This refinement brackets the minimizer and provides the required numerical precision.
The minimum is sandwiched between $3.5492598$ and $3.54925986$. For $\tau_{0}=1.6469768608776936$ (exact value) we obtain $\xi=0.8119098734258519\ldots$, $\theta=0.5909025598581181\ldots$, and a corresponding inspective curve with minimum distance at least $0.0302318\ldots$ to the disk boundary (corresponding to $\tau(\chi)\approx 0.24774522\ldots$ for some $\chi$), and reported cost approximately $3.5492595860809693\ldots$. Numerical accuracy of this computation is discussed in Section 5.4. ∎
@@ 5.4 Numerical Methods and Accuracy Guarantees
We used the Julia programming language [bezanson2017julia] for all computations.
For the convex nonlinear program that lower bounds $s(\theta)$ in Lemma 5.9, we model it with JuMP [dunning2017jump] and solve it using the interior point method Ipopt [Ipopt, wachter2006implementation]. The program is convex since the feasible set $\{t_{i}\geq 0\}$ is convex and the objective is a nonnegative conical combination of norms of affine maps. Hence any KKT point is globally optimal. In practice, the solver returns primal and dual feasible solutions with residuals below $10^{-14}$ and objective values stable to at least $10^{-9}$. Because convexity guarantees global optimality, these certificates validate the solutions and support the digits reported in our lower bounds.
For the ODE system of Definition 2.5, used in Section 5.3 to establish feasibility of trajectories and to evaluate the cost functional, we proceed as follows. The equation for $\psi$ has a singularity at $x=0$, so we begin the integration at $x_{0}=10^{-6}$ using the asymptotic expansion $\psi(x_{0})=\pi/2-\pi x_{0}+(\pi^{2}/2)x_{0}^{2}$, whose truncation error is $O(x_{0}^{3})\approx 10^{-18}$ and therefore negligible compared with the solver tolerances. We integrate $\psi$ on $x_{0},1]$ with the adaptive stiff solver Rodas5() from the DifferentialEquations.jl library [rackauckas2017differential], setting absolute and relative tolerances to $10^{-12}$, and then integrate $\tau$ on the same interval using the dense output of $\psi$ in the right-hand side.
The deployment parameter $\xi$ is obtained in two stages. A uniform grid of 10,00010{,}000 points provides a coarse $\xi_{\text{approx}}$ as the last feasible point according to a geometric predicate that tolerates floating-point noise at the level $10^{-12}$. This value is refined by bisection to absolute tolerance $10^{-8}$. To further confirm stability, we repeat the bisection at tolerance $5\cdot 10^{-10}$ and recompute the objective, reporting the discrepancy between the two runs, which is consistently negligible. In addition, the minimum of $\tau(\cdot)$ on $x_{0},\xi_{\text{approx}}]$ is computed by Brent’s method [brent2013algorithms] with tolerance $10^{-10}$ to certify that all trajectories remain uniformly outside the disk.
The integral of Lemma 5.4 is evaluated using the adaptive Gauss-Kronrod quadrature routine quadgk from QuadGK.jl [quadgk], with relative tolerance $10^{-12}$ and absolute tolerance $10^{-14}$. Substituting $(\xi,\theta)$ into the expression of Theorem 2.8 then yields the values reported in Section 5.3.